OpenStudy (anonymous):

Any help would be great!! graph the inequality 4x - 2y < -3

OpenStudy (callisto):

First, graph the equation 4x - 2y = -3 Pick a testing point (0, 0) LS = 4(0) - 2(0) = 0 > -3 So, (0, 0) is NOT included in the shaded region. In other words, shade the region where (0, 0) is not located at.

OpenStudy (anonymous):

that helps...but i don't know how to graph 4x - 2y = -3

OpenStudy (callisto):

Put x =0 into the equation to find the corresponding y-value. (0, y) Put y =0 into the equation to find the corresponding x-value. (x, 0) Plot these 2 points on a graph with marked scale. Draw a line joining the 2 points. Write the equation next to the line.

OpenStudy (anonymous):

 so would the end graph look something like this?

OpenStudy (callisto):

Yes, I think!

OpenStudy (anonymous):

great thanks so much for you help!

OpenStudy (callisto):

Welcome!